OpenStudy (anonymous):

If h(x) is the inverse function of f(x), find h'(5)

OpenStudy (anonymous):

\[f\circ h(x)=x\\ f'(h(x))\times h'(x)=1\] by the chain rule

OpenStudy (anonymous):

that means \[h'(x)=\frac{1}{f'(h(x))}\]

OpenStudy (anonymous):

of course that doesn't tell you what \(h'(5)\) is unless you know \(f(x)\)

OpenStudy (anonymous):

it just gives me a table of values for each function f(x) 5,8; f'(x) 4, 6; f"(x) -3,10

OpenStudy (anonymous):

well maybe that is enough the numbers you need are \(h(5)\) and \(f'\) of that number

OpenStudy (anonymous):

can you post a screen shot or something? this is a bit hard to read

OpenStudy (anonymous):

yes just a second

OpenStudy (anonymous):

ok i see it from the table \[f(3)=5\]right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

that means \(h(5)=3\) so now we are at \[\frac{1}{f'(h(5))}=\frac{1}{f'(3)}\]

OpenStudy (anonymous):

if you can find \(f'(3)\) in the table you are done

OpenStudy (anonymous):

no those are the only numbers given so i only see f'(4)

OpenStudy (anonymous):

hmm let me look again

OpenStudy (anonymous):

oh no, you are definitely told what \(f'(3)\) is look at the bottom row of the table

OpenStudy (anonymous):

5?

OpenStudy (anonymous):

is it not clear how to read the table?

OpenStudy (anonymous):

oh. I see now. Thank you so much for your help!

OpenStudy (anonymous):

oh wow sorry you are right, i was reading it wrong !

OpenStudy (anonymous):

the top row has a 3, and so yes, \(f'(3)=4\)

OpenStudy (anonymous):

that makes your answer \(\frac{1}{4}\)

OpenStudy (anonymous):

got it thank you so much! any chance you have time for one more? I can take a screen shot of it as well

OpenStudy (anonymous):

yes, a screen shot will help

OpenStudy (anonymous):

#15. the one that is circled

OpenStudy (anonymous):

you need the derivative of arctangent first do you know it?

OpenStudy (anonymous):

if not you can look it up

OpenStudy (anonymous):

\[\frac{ 1 }{ 1+x ^{2}? }\]

OpenStudy (anonymous):

yes, and so \[f'(\sqrt{3})=\frac{1}{1+\sqrt{3}^2}=\frac{1}{4}\]

OpenStudy (anonymous):

that is the slope of your line now you need the point on the graph of \(f(x)=\arctan(x)\) and \(x=\sqrt3\) do you know what \(\arctan(\sqrt3)\) is ?

OpenStudy (anonymous):

if not, i will tell you

OpenStudy (anonymous):

.939?

OpenStudy (anonymous):

oh my you need an exact answer, not a decimal it is \(\frac{\pi}{3}\)

OpenStudy (anonymous):

i just typed it in my calculator i didn't know it was supposed to be in radians

OpenStudy (anonymous):

i just pulled up a unit circle and I see it now

OpenStudy (anonymous):

yeah of course in any case you can find the equation of the line with slope \(\frac{1}{4}\) through the point \((\sqrt3,\frac{\pi}{3})\)

OpenStudy (anonymous):

use the point slope formula \[y-\frac{\pi}{3}=\frac{1}{4}(x-\sqrt3)\]

OpenStudy (anonymous):

right, so is k 3root3/12 plus 4pie/12?

OpenStudy (anonymous):

i guess, i didn't do it but it is something ugly like that

OpenStudy (anonymous):

ok i just wanted to make sure i wasn't supposed to compute that out into one single number with a decimal

OpenStudy (anonymous):

no, i it is whatever you get when you write the line in standard form multiply by 4 etc

OpenStudy (anonymous):

ok great thank you. you were a lot of help

OpenStudy (anonymous):

yw